OpenStudy (anonymous):

Find exact solutions |4y-3|=5.

jigglypuff314 (jigglypuff314):

|4y - 3| = 5 can be separated into two equations: 4y - 3 = 5 and 4y - 3 = -5 solve for y for both to do that, add 3 to both sides then divide both sides by 4 you should get two answers :)